OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y = x + 5 and that pass through the point (-2, -1). parallel: y = -x - 1 perpendicular: y = x + 2 parallel: y = x - 1 perpendicular: y = -x + 1 parallel: y = x + 1 perpendicular: y = -x - 3 parallel: y = 2x - 2 perpendicular: y = -2x - 1

OpenStudy (anonymous):

ok, start by determine what the slope of y=x +5

OpenStudy (anonymous):

Ok, but I'm not really sure of how to solve for it in general.

OpenStudy (anonymous):

x-y+5=0 slope=-(coefficient of x)/(coefficient of y)

OpenStudy (anonymous):

parallel: y = -x - 1 perpendicular: y = x + 2 parallel: y = x - 1 perpendicular: y = -x + 1 parallel: y = x + 1 perpendicular: y = -x - 3 parallel: y = 2x - 2 perpendicular: y = -2x - 1 Those were the answer choices

OpenStudy (anonymous):

i know these options ........... u just tell me the slope

OpenStudy (anonymous):

Hint: parallel lines have same slopes the product of slopes of two perpendicular lines =-1

OpenStudy (anonymous):

y= mx+b @nitz you dont need to change into standard form m= slope so for y = x + 5 m=1 for parallel lines, the slope is the same for perpindicular lines, the slope of the perpendicular line is the negative recipocal so once you have these two slopes, substitute them with the point into the point slope form and then simplify to slope intercept form \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

Oh ok, thank you. :)

OpenStudy (campbell_st):

A simple way to do this problem is to just graph the pairs of lines... try geogebra http://www.geogebra.org/cms/ a great free graphing geometry download